OpenStudy (anonymous):

What is the average binding energy per nucleon for a C-12 nucleus with a mass defect of 0.0993 amu? (1 amu= 1.66 x 10-27 kg; 1 J = 1 kg m2/s2)

OpenStudy (rogue):

You can find average binding energy via this equation:\[E = \Delta m c^2 =(Zmp + Nmn - Ma) \times 931.494 MeV/u\]Z = atomic # mp = mass of proton N = # neutrons mn = mass of neutron Ma - atomic amss